OpenStudy (dranz10):

Solve 2 over 3 (x − 4) = 2x . I'm having trouble knowing what to do with the fraction.... I'll medal. Thanks!

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

if you hate the fraction, multiply both sides by 3 and start with \[2(x-4)=6x\]

OpenStudy (dranz10):

If I multiplied by three why would it be 2(x-4) and not 3(x-4)

OpenStudy (dranz10):

@misty1212

OpenStudy (michele_laino):

since your equation is like this: \[\frac{2}{3} \cdot \left( {x - 4} \right) = 2x\]

OpenStudy (michele_laino):

If I multiply both sides by \(3\), I get: \[\frac{2}{3} \cdot 3 \cdot \left( {x - 4} \right) = 2x \cdot 3\]

OpenStudy (dranz10):

uhu

OpenStudy (michele_laino):

so, after a simplification, we can write: \[2\left( {x - 4} \right) = 6x\]

OpenStudy (dranz10):

Oh.....I get it now..Thank you!!

OpenStudy (michele_laino):

:)